Nitrogen and hydrogen gases are present in a rigid steel container which react to form ammonia gas (NH3). Use the information below to answer the following questions.

N2(g) + 3H2(g) --> 2NH3(g)

 Before this chemical reaction took place, about 98.0 grams of nitrogen gas were present along with some mass of hydrogen gas.

 After the chemical reaction has occurred, 51.0 g ammonia plus 56.0 grams of excess reactant (either nitrogen or hydrogen) are present in a rigid container.

a. Identify the limiting reactant, or state that neither reactant is limiting. Circle your selection below and then justify your answer. Limiting: Nitrogen Hydrogen Neither

b. Determine the mass of hydrogen that was present in the container before the reaction. Show work which clearly displays all steps.
Answer #1

Moles of N2 = 98/28 = 3.5 moles

N2 + 3H2 -----> 2NH3

Moles of NH3 produced = 51/17 = 3 moles

Limiting reagent will drive the yield of NH3.

From reaction;

1 mole N2 produces 2 moles NH3

So, 3.5 mole will produce = 2*3.5 = 7 moles NH3

However, available moles of NH3 = 3 which is less than 7 moles.

Ideally, if N2 was limiting reagent so it would produce 7 moles of NH3. Thus, N2 is excess reagent and H2 is limiting reagent.

Excess moles of N2 = 56/28 = 2 moles

Used moles of N2 = 3.5 - 2 = 1.5 moles

From reaction;

1 mole N2 requires 3 moles H2

So, 1.5 moles will require = 1.5 * 3 = 4.5 moles of H2

Mass of H2 = moles * molar mass = 4.5 * 2 = 9 grams ....Answer